2. Let Qd stand for the quantity demanded, Qs stand for the quantity supplied, and P stand for price. If Qd = 20 - 2P and Qs = 5 + 3P, then the equilibrium price is


1
Expert's answer
2021-06-11T11:23:30-0400

Equilibrium price is calculated at a point where Quantity demanded is equal to quantity supplied.

"Qd=Qs\\20-2p=5+3p\\20-5=3p+2p\\15=5p\\p=3"

Therefore equilibrium price is 3